The height h(t), in feet, of an airborn tee-shirt t seconds after being launched can be approximated by h(t)=-15t^2+105t+10, 0<= t <=10. Find the time when the shirt will reach a fan 160ft above ground level?
Answer by nerdybill(7384)

The height h(t), in feet, of an airborn tee-shirt t seconds after being launched can be approximated by h(t)=-15t^2+105t+10, 0<= t <=10. Find the time when the shirt will reach a fan 160ft above ground level?
.
set h(t) to 160 and solve for t:
h(t)=-15t^2+105t+10
160=-15t^2+105t+10
0=-15t^2+105t-150
0=15t^2-105t+150
0=t^2-7t+10
0=(t-2)(t-5)
t = {2, 5}
.
answer: at 2 secs and again at 5 secs